### v2

# REIC: RAG-Enhanced Intent Classification at Scale

Abstract
Accurate intent classification is critical for efficient routing in customer service, ensuring customers are connected with the most suitable agents while reducing handling times and operational costs. However, as companies expand their product lines, intent classification faces scalability challenges due to the increasing number of intents and variations in taxonomy across different verticals. In this paper, we introduce REIC, a Retrieval-augmented generation Enhanced Intent Classification approach, which addresses these challenges effectively. REIC leverages retrieval-augmented generation (RAG) to dynamically incorporate relevant knowledge, enabling precise classification without the need for frequent retraining. Through extensive experiments on real-world datasets, we demonstrate that REIC outperforms traditional fine-tuning, zero-shot, and few-shot methods in large-scale customer service settings. Our results highlight its effectiveness in both in-domain and out-of-domain scenarios, demonstrating its potential for real-world deployment in adaptive and large-scale intent classification systems.
REIC: RAG-Enhanced Intent Classification at Scale
Ziji Zhang, Michael Yang, Zhiyu Chen, Yingying Zhuang, Shu-Ting Pi, Qun Liu, Rajashekar Maragoud, Vy Nguyen, Anurag Beniwal Amazon.com Inc, Seattle, USA {czhangzi,abyang,zhiyuche,yyzhuang,shutingp}@amazon.com {qunliu,maragoud,nguynvy,beanurag}@amazon.com
1 Introduction
Customer service cui2017superagent; chen-etal-2024-identifying; nguyen-etal-2020-dynamic; qi-etal-2021-benchmarking; chen-etal-2023-generate; zhou-etal-2023-towards-open; pi2024topology is critical for modern e-commerce but also one of the most resource-intensive departments. Different agents, either human or model, are trained to handle specific types of customer issues, making precise intent classification, particularly at the issue level, crucial for efficient routing. High issue-oriented intent accuracy ensures that customers are connected with the most suitable agents, reducing unnecessary transfers and lowering handling times. This optimization not only enhances customer satisfaction but also cuts operational costs by streamlining interactions and improving overall service efficiency. For model-based automatic resolvers in chatbot agentic systems gupta2024dard, the ability to precisely identify user intent is essential for delivering contextually appropriate and solution-oriented responses.
As companies expand their product lines, intent classification faces two key challenges. First, the number of customer intents grows over time, requiring models to adapt to new intents quickly. Second, intent taxonomies can vary across product lines, making it difficult to maintain a unified classification systempi2024universal. For example we organize products into different verticals in e-commerce: with third-party products, customer usually inquire about physical retail orders or consumer accounts, and intents are categorized into three levels from coarse to fine-grained. In contrast, first-party products require more customized customer services due to our proprietary device and digital product offerings. For instance, a customer seeking device troubleshooting may interact with an agent who can access real-time diagnostic information and perform specific troubleshooting steps on the user’s behalf. This necessitates a broader set of intent categories to accommodate diverse customer needs, as illustrated in Figure 1. This heterogeneity complicates intent classification, demanding scalable and flexible approaches to ensure accurate routing and efficient customer service. In this work we demonstrate our method using two verticals but it can be easily adapted to more.
In this paper, we propose a novel Retrieval-augmented generation Enhanced Intent Classification (REIC) approach that reduces computational complexity and improves scalability for intent classification. We demonstrate the effectiveness of this approach through extensive experiments on real-world datasets , showing that our method outperforms traditional fine tuning and zero-shot or few-shot methods in large-scale customer service settings. Our results on both in-domain and out-of-domain intents demonstrate its potential to improve classification accuracy and enable dynamic updates without retraining, making it ideal for industry-scale applications.
Figure 1: We present the heterogeneous intent structure with representative examples, illustrating the intent label hierarchy in each vertical.
2 Related Work
2.1 Intent classification
Early work on intent classification for dialogues often relied on bag-of-words or recurrent models. For example, schuurmans2019intent evaluated various classifiers on a multi-domain intent dataset and found that a simple SVM with hierarchical label taxonomy outperformed deeper LSTM models. With the advance of transformer architectures, researchers began to leverage self-attention and multi-task learning for intent understanding. ahmadvand2020jointmap introduced a joint intent mapping model that simultaneously classifies high-level intent and maps queries to fine-grained product categories. wang2021mell employed a slowly updated text encoder and global/local memory networks to mitigate catastrophic forgetting and parameter explosion for large-scale intent detection task. Recent work has pushed toward using large pre-trained models and retrieval-based prompting to enable cross-domain and zero/few-shot intent classification. liu2024lara proposed a framework which integrates a fine-tuned XLM-based intent classifier with an LLM to essentially treat multi-turn intent understanding as a zero-shot task. yu2021few also explored retrieval-based methods for intent classification and slot filling tasks in few-shot settings. Our work adopts similar in-context learning (ICL) setup while focusing on handling large-scale multi-domain intent classification task from industry level applications.
2.2 In-context Learning
The performance of LLM has been significantly enhanced in few-shot and zero-shot NLP tasks through ICL. Recent ICL research focus on how to effectively identify and interpret retrieved context. guu2020realmretrievalaugmentedlanguagemodel first showed how to pre-train masked language models with a knowledge retriever in an unsupervised manner. karpukhin2020densepassageretrievalopendomain proposed a training pipeline in which retrieval is implemented using dense representations alone and embeddings are learned from a small number of questions and passages with a dual-encoder. ram2023incontextretrievalaugmentedlanguagemodels considered simple alternatives to only prepend retrieved grounding documents to the input, instead of modifying the LLM architecture to incorporate external information. Similar approaches have proven particularly effective in the application of RAG on dialogue systems shuster2022blenderbot3deployedconversational; shuster-etal-2022-language, specifically goal-oriented and domain-specific dialogs from customer service scenarios zhuang-etal-2021-weakly; 10.1145/3543873.3587680. In our work, we utilize ICL in both LLM fine-tuning stage for data generation and at inference-time with an intent candidate retriever.
Figure 2: The proposed REIC method from customer query to routing intent leveraging on vector retrieval and probability calculation.
3 Preliminary
Intent classification for queries is typically framed as a multiclass text classification problem. Specifically, given a customer query $q\in Q$, the goal is to map it to one of the $k$ pre-defined intents $t\in T=\{t_{1},...,t_{k}\}$ using a model $\mathcal{M}$ so that the predicted intent $\hat{t}=\mathcal{M}(q)$ maximizes the probability of correctly classifying $q$. Formally, this can be expressed as:
$\hat{t}=\arg\max_{t\in T}\,P(t\mid q;\theta)$ (1)
where $P(t\mid q;\theta)$ denotes the probability of intent $t$ given query $q$, parameterized by $\theta$ of the model $\mathcal{M}$.
However, intent labels are often not mutually independent; correlations between intents can exist, making a flat classification structure suboptimal. This leads to two major challenges in large-scale, industry-level intent classification: 1) scalability: a large number of intent labels $k$ can make flat classification computationally expensive and difficult to scale, especially as the set of intents grows; 2) label correlation: related intents, such as "Order Issue" and its sub-intents “Track Order” or “Cancel Order”, are treated independently in flat classification, ignoring their hierarchical relationships and increasing the risk of misclassification.
To address these issues, hierarchical intent classification is preferred. In this approach, a query $q$ is classified progressively from general categories to specific sub-intents, enhancing both efficiency and accuracy for industry-scale applications.
Note that for a more generalized setting, intents from different verticals and domains may have entirely different hierarchy and ontology. In Figure 1, we demonstrated some examples of intent hierarchy in our application, which involves customer service query intent detection with two business verticals: Third-Party or 3P business (customer contacting about third-party physical retail orders or consumer accounts) and First-Party or 1P business (customer contacting about first-party digital or device issues). Both verticals span a diverse range of product types, reflecting the broad scope of customer inquiries handled by our system. If we use the traditional single-head flattened intent labels, the total intent ontology set size would be at $10^{3}$ level which create major challenges for accurate intent classification. By creating hierarchical intent ontology across different business verticals, each classification head only needs to handle less than 50 intents that are more manageable for language models. In the following sections of this paper, we utilize this intent ontology setup for experiments and comparisons.
4 Method
LLMs has revolutionized the landscape of customer engagement, particularly in the domain of intent detection systems. While these models demonstrate remarkable capabilities in language comprehension and knowledge representation, their adaptation to domain-specific contexts presents notable challenges. Specifically, the integration of industry-specific terminology, organizational nomenclature, and distinctive customer service scenarios necessitates fine-tuning and customization of these models. To address these limitations and enhance the accuracy of customer intent identification, we introduce a novel approach REIC for RAG-Enhanced Intent Classification.
REIC aims to bridge the gap between the generalized capabilities of LLMs and the specialized requirements of diverse business environments. By leveraging RAG, our approach seeks to enhance the precision and relevance of intent detection, thereby facilitating more nuanced and context-aware customer interactions across various industry-specific scenarios. The method consists of three main components: index construction, candidate retrieval, and intent probability calculation (Figure 2).
Index Construction
We first construct a dense vector index containing (query, intent) pairs from a held-out annotated dataset. Each query is encoded using a pre-trained sentence transformer model to generate dense vector representations. The corresponding intent labels are stored alongside these embeddings. The intent labels follow a hierarchical structure with $d$ dimensions which represents different intent domain knowledge and might range from different domains.
Candidate Retrieval
Given a new query $q$, we first encode it using the same encoder for index construction to obtain its dense vector representation $\mathbf{v}_{q}$. We then perform approximate nearest neighbor search to retrieve the top-k most similar (query, intent) pairs, denoted as set $E$. The similarity is computed using cosine distance between the query vector and indexed vectors:
$sim(q,q_{i})=\frac{\mathbf{v}_{q}\cdot\mathbf{v}_{i}}{\|\mathbf{v}_{q}\|\|\mathbf{v}_{i}\|}$ (2)
where $\mathbf{v}_{i}$ represents the vector encoding of the i-th indexed query.
Intent Probability
For the retrieved set $E$, we leverage a fine-tuned LLM $\mathcal{M}$ to perform constrained decoding and calculate the probabilities of the possible intents. Given a prompt template $\mathcal{P}$, the LLM takes as input the instantiated prompt, which includes the original query $q$ and the retrieved (query, intent) pairs as context. For each unique intent $t_{j}$ in $E$, we compute:
$P(t_{j}|q,E)=\mathcal{M}(\mathcal{P},q,E)_{t_{j}}$ (3)
where $\mathcal{M}(\mathcal{P},q,E)_{t_{j}}$ represents the model’s predicted probability for $t_{j}$ given the query and retrieved examples.
The final intent classification is determined by selecting the intent with the highest probability:
$\hat{t}=\operatorname*{arg\,max}_{t_{j}\in E}P(t_{j}|q,E)$ (4)
This approach enables dynamic updates to the intent space by simply adding new (query, intent) pairs to the index, leveraging the in-context learning capabilities of the LLM without requiring model retraining.
Figure 3: Constrained decoding for probability calculation.
The probability-based reranking helps mitigate potential LLM hallucination by grounding predictions in retrieved examples (Figure 3). With tradition greedy decoding, sometimes the LLM might generate intents outside of the given candidate list and cause downstream routing failure. We perform constrained decoding to calculate the probability of each retrieved intent $t_{j}$ in $E$, which ensures the success of downstream routing. Given prompt $\mathcal{P}$ with instructions, retrieved candidates $E$, and customer query $q$, we append $t_{j}$ at the end to calculate the total logits from model forward pass $\mathcal{L}_{t_{j}}=\mathcal{M}(\mathcal{P}(E,q)+t_{j})$. Then we mask out the positions of $\mathcal{P}(E,q)$ and accumulate the log probabilities for the intent sequence $t_{j}$ with length $s_{j}$:
$\mathcal{M}(\mathcal{P},q,E)_{t_{j}}=\exp{(\sum_{t_{j}}\text{LogSoftmax}(\mathcal{L}_{t_{j}})/s_{j})}$ (5)
During training, we train the intent LLM $\mathcal{M}$ by minimizing the cross-entropy loss between the predicted and ground-truth intents. During inference, instead of traditional auto-regressive next token decoding, we perform one model forward-pass calculation with a batch size $k$ for top-k intent candidates and get the $k$ probabilities for re-ranking and final intent prediction.
5 Experimental Setup
5.1 Datasets
Due to business considerations, we are not permitted to share the results using the original customer data. As a result, we manually anonymized both the labels and transcripts to ensure no personal information is included. Additionally, specific product and service names were denonymized to prevent the identification of the company from the transcript or label descriptions. Despite these modifications, the conclusions drawn from our experiments remain valid. The final dataset contains 52,499 training samples with 35,041 1P Business queries and 17,458 3P Business queries. The test set consists of 3,647 1P Business queries and 1,717 3P Business queries respectively using random sampling. All of the data samples have incorporated retrieved intent candidates from the retriever. We also performed dataset cleaning in the training set to make sure the true intent is contained in the retrieved list. During inference, we use the actual noisy retrieved list which also relies on the capability of the embedding model.
Models 3P Business vertical 1P Business vertical Overall
Precision Recall F1 Precision Recall F1 Precision Recall F1
RoBERTa 0.527 0.447 0.483 0.583 0.488 0.531 0.565 0.474 0.516
Mistral Classification 0.215 0.228 0.221 0.301 0.250 0.273 0.269 0.243 0.255
Claude Zero-shot 0.338 0.250 0.287 0.238 0.170 0.199 0.271 0.196 0.227
Claude Few-shot 0.386 0.289 0.331 0.350 0.308 0.328 0.361 0.302 0.329
Claude + RAG 0.473 0.438 0.455 0.415 0.389 0.402 0.434 0.404 0.419
REIC 0.538 0.546 0.542 0.600 0.574 0.587 0.579 0.565 0.572
Table 1: Intent detection confusion matrix on different business with different methods
5.2 Compared Methods
We consider the following baselines:
  • RoBERTa: We fine-tune RoBERTa-base liu2019roberta with multiple classification heads. This adaptation allowed the model to simultaneously categorize utterances across multiple dimensions.
  • Mistral Classification [^1] Due to legal concerns, we are not permitted to use non-commercial LLMs like Llama.: We fine-tune a Mistral-7B-v0.3222https://huggingface.co/mistralai/Mistral-7B-v0.3 with a sequence classification head. Instead of directly generating output sequences, the model projects the pooled embedding into a space with the same dimension as the number of classes.
  • Claude Zero-shot: We employ the Claude 3.5 Sonnet model in a zero-shot configuration. To facilitate accurate intent prediction, we craft a comprehensive prompt that explicitly defines each potential intent.
  • Claude Few-shot: Similar to Claude Zero-shot, we incorporate 20 demonstration examples, with 10 from each vertical, to enhance coverage of diverse intents across different domains.
  • Claude+RAG: Instead of using a fine-tuned LLM, we employ Claude 3.5 Sonnet as the backbone and incorporate the same set of retrieved candidates as described in Section 4 into the prompt. This comparison allows us to assess whether a smaller fine-tuned LLM can perform competitively against a large foundation model for this task.
5.3 Implementation Details
The LLM component of our REIC approach utilizes a fine-tuned model from Mistral-7B-Instruct-v0.2333https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2. We applied 8 NVIDIA-A100 40GB GPUs with 96 vCPUs to conduct PEFT peft training with LoRA adapters hu2022lora. We choose a set of LoRA parameters with a rank of 8, an alpha value of 16, and a dropout rate of 0.1. The training batch size is set to 8 per GPU with a learning rate of $2e^{-5}$. We train the model using Cross Entropy Loss for 3 epochs which takes around 3 hours on the instance.
We experimented with the following off-the-shell retrievers for candidate retrieval:
  • BM25 robertson1995okapi is a widely used traditional sparse retrieval method. Although it is unsupervised, it consistently demonstrates strong performance across a variety of benchmarksthakur2021beir.
  • MPNet [^4] https://huggingface.co/sentence-transformers/all-mpnet-base-v2 song2020mpnet is a sentence embedding model fine-tuned on one billion sentence pairs using a contrastive learning objective.
  • Contriever-MS MACRO izacard2022unsupervised is an unsupervised dense retriever pre-trained with contrastive learning and fine-tuned on MS MARCO nguyen2016ms.
  • ColBERT-v2 santhanam-etal-2022-colbertv2 is a late-interaction retriever that combines denoised supervision and residual compression to improve retrieval quality and reduce space footprint.
6 Results
6.1 Intent Detection Ability
To evaluate the effectiveness of our REIC method in intent detection, we conducted experiments comparing it against several baseline methods described in §5.2. Our results, presented in Table 1, illustrate performance across two business verticals (3P Business and 1P Business) and an overall aggregate assessment based on Precision, Recall, and F1-score.
The results indicate that our REIC offers significant advantages in intent detection over standard fine-tuning or prompting-based methods. While fine-tuned models like RoBERTa perform reasonably well, they require extensive retraining when new intents emerge. We hypothesize that the limited performance of the Mistral Classification model stems from its nature as a decoder-only architecture, which may be less effective in extracting the semantic meaning of input query. Additionally, since it is not pretrained for classification tasks, incorporating a classification head during fine-tuning is unlikely to yield optimal results. Prompting-based approaches (Claude Zero-shot and Few-shot) generally underperformed, with Claude Few-shot achieving a maximum F1-score of 0.329 overall. The Claude + RAG method improved performance compared to standalone prompting but remained inferior to our approach by 26.7%.
These observations confirm that the integration of RAG and fine-tuned LLM enables greater flexibility, improved precision, and higher recall rates, making it well-suited for handling diverse and evolving intent spaces in different applications.
6.2 Impact of Retrievers
In order to evaluate the impact of retrievers on the final performance, we experimented four different retrievers in REIC including one sparse retrieval method and three dense retrievers, details in §5.3. The intent detection accuracy across different business verticals using these retrievers is presented in Table 2.
BM25, despite being an unsupervised sparse retrieval method, performs competitively, achieving an overall accuracy of 0.532. Among the dense retrievers, MPNet outperforms the others, attaining the highest accuracy across both the 3P Business and 1P Business verticals. This suggests that MPNet’s contrastive learning-based sentence embeddings are highly effective for retrieving relevant candidates that aid intent classification. In contrast, Contriever exhibits the lowest accuracy across all categories.
Our findings show that retriever selection significantly impacts intent classification. Although BM25 is a strong baseline, dense retrievers like MPNet consistently outperform it. This highlights the value of high-quality embeddings and extensive fine-tuning on large datasets, which is why we have chosen MPNet as our final retriever in REIC.
Retriever 3P Business 1P Business Overall
BM25 0.521 0.537 0.532
Contriever 0.450 0.461 0.457
ColBERTv2 0.503 0.560 0.542
MPNet 0.545 0.573 0.564
Table 2: Intent detection accuracy on different business verticals using different retrievers in REIC
7 Impact of Retrieval Candidate Size
We investigated the impact of different retrieval candidate numbers (top-k) in REIC to balance intent detection accuracy and inference latency. The Figure 4 illustrates the trade-off between these two factors, with overall accuracy plotted on the left y-axis (blue) and inference latency on the right y-axis (red) against different values of top-k. From the accuracy perspective, increasing top-k allows the model to access a broader range of relevant information, leading to better predictions. Beyond a certain threshold, additional retrieved candidates contribute minimally to accuracy while still increasing computational complexity. Latency, on the other hand, exhibits a sharp rise as top-k increases. This indicates a crucial trade-off: although retrieving more candidates can improve accuracy, it also leads to longer inference times, which may not be suitable for real-time applications.
In our experiments, we select top-k = 10 which ensures a meaningful accuracy boost without incurring excessive computational costs. However, the ideal top-k may vary depending on application requirements. For instance, real-time systems such as customer service chatbots or voice assistants may favor a lower top-k to maintain fast response times. Conversely, offline or batch-processing applications could accommodate higher top-k values if maximizing accuracy is a priority. Our findings emphasize the need to carefully tune retrieval parameters in REIC to meet specific operational demands.
Figure 4: The accuracy and latency when using different retrieval top-k values.
7.1 Robustness on Unseen Intents
To evaluate REIC’s robustness on unseen intents, we trained our models exclusively on the 3P Business vertical and tested them on the 1P Business vertical, simulating a real-world out-of-domain scenario. As illustrated in Figure 1, 1P Business vertical has 4 intent category with more than 800 unique intent combinations, while the training data used from 3P Business vertical has 3 intent category with only around 70 unique intents. This out-of-domain scenario helps assess how well REIC generalizes to new, previously unseen intents. The results are summarized in Table 3.
Claude Zero-shot performs the worst, with an accuracy of 0.17 on the 1P Business vertical. In contrast, Claude Few-shot shows improvement, achieving an accuracy of 0.308 on the 1P Business vertical. This demonstrates that providing a few examples significantly enhances the model’s ability to generalize.
Notably, the RAG-based methods, particularly Claude + RAG, significantly outperform Claude Few-shot, achieving 0.389 on the 1P Business vertical. This demonstrates the advantage of our RAG-based strategy in handling unseen intents, as it dynamically retrieves the most relevant examples to enhance predictions, surpassing static few-shot examples. Similarly, REIC, although slightly lower than Claude + RAG, still performs strongly compared to Claude Few-shot, highlighting the model’s effectiveness on both in-domain and out-of-domain intents. Overall, REIC excels in-domain, and its performance on the 1P Business vertical remains competitive with Claude Few-shot, underscoring the robustness and adaptability of our REIC approach for unseen domains.
Models 3P Business 1P Business Overall
Claude Zero-shot 0.250 0.170 0.196
Claude Few-shot 0.289 0.308 0.302
Claude + RAG 0.438 0.389 0.404
REIC 0.538 0.283 0.364
Table 3: Out-of-domain intent detection accuracy
7.2 Online Deployment Performance
Following the deployment of our REIC in an internal system, we observed significant improvements in online performance, particularly in intent detection accuracy. Compared to the previously deployed system, which relies on two separate models using traditional fine-tuning approaches for different business verticals, our REIC method simplified the system using just one consolidated model, which reduced misclassifications and improved resolution routing. To measure the success rate, we include a confirmation question for the customer to verify whether the predicted intent or issue is correct. Our REIC leads to a 3.38% absolute improvement for customer positive response rate. These improvements validate the efficacy of REIC for real-world deployment, offering both accuracy gains and operational efficiency in intent classification.
8 Conclusion
This paper presents a novel RAG-Enhanced Intent Classification (REIC) method that addresses scalability challenges and the heterogeneity of intent taxonomies in large-scale customer service systems. By incorporating a hierarchical intent classification strategy, REIC significantly reduces computational complexity. Leveraging the RAG technique, our method dynamically integrates contextually relevant retrieved examples, outperforming traditional fine-tuning, as well as zero-shot and few-shot approaches, in intent detection tasks. Additionally, our results demonstrate strong performance on both in-domain and out-of-domain test sets, highlighting its applicability for industry-scale applications.
9 Limitations
While REIC demonstrates strong performance in both in-domain and out-of-domain intent classification, it has a few limitations. Its accuracy remains closely tied to the quality of the retriever: if the correct intent is not among the retrieved candidates, the model cannot recover, underscoring the need for more robust retrieval methods or fallback mechanisms. In addition, although REIC allows for dynamic updates without retraining, it still relies on a fixed number of retrieved candidates, creating a trade-off between accuracy and latency that may hinder its deployment in real-time applications. Future work could address these challenges by developing adaptive retrieval strategies or introducing confidence-based mechanisms to dynamically adjust the candidate pool.